When asked about her age, Jun's blind date answered: "the sum of the 2 digits of my age is 8. 14 years from now, mty age will be 4 less han my present age in reverse" How old is Jun's date?
Answer by ankor@dixie-net.com(22740)

"the sum of the 2 digits of my age is 8."
14 years from now, my age will be 4 less than my present age in reverse"
How old ?
:
let a = the first digit of his age
let b = the 2nd digit
then
10a+b = his age
and
10b+a = the reversed age
:
Write an equation, for each statement:
:
"the sum of the 2 digits of my age is 8."
a + b = 8
:
14 years from now, my age will be 4 less than my present age in reverse"
10a+b + 14 = 10b+a - 4
10a - a = 10b - b - 4 - 14
9a = 9b - 18
simplify, divide by 9
a = b - 2
a - b = -2
add to the 1st equation
a - b = -2
a + b = 8
-------------Adding eliminates b, find a
2a = 6
a = 3 is the first digit of his age
then obviously
b = 5 is the 2nd digit
:
He is 35 yrs old
:
:
:
Check this using these values in the original equation
10a+b + 14 = 10b+a - 4
35 + 14 = 53 - 4